The concentration of zinc is 0.5M and copper is 0.05M at 25 0C. Using the Nernst equation calculate the cell potential of Zn/Zn2+//Cu2+/Cu.
1
Expert's answer
2020-12-08T06:59:38-0500

The concentration of zinc is 0.5M and copper is 0.05M at 25 0C. Using the Nernst equation calculate the cell potential of Zn/Zn2+//Cu2+/Cu.

The equations of reactions taking place in the cell are the following:

Cu2+ + 2 e ⇄ Cu(s)     E0Cu2+/Cu = 0,34 V: cathode

Zn(s) ⇄ Zn2+ + 2 e      E0Zn2+/Zn = -0,76 V: anode

Cu2+(aq) + Zn(s) → Cu(s) + Zn2+(aq).

The Nernst equation for the potential of the cell is:

"E = E^0 - \frac{RT}{nF}\text{ln}Q" , where:

  • n is the number of the electrons taking part in the reaction, 2 in our case
  • R = 8,314 J.mol–1.K–1 : ideal gas constant
  • T : absolute temperature in Kelvin
  • F = NA × e = 6.022 141 79x1023 × 1.602 176 487x10–19 = 96485.3399 C : Faraday constant (charge of 1 mole of electrons)
  • "Q" is the reaction quotient, "\frac{c(Zn^{2+})}{c(Cu^{2+})}" .


Therefore, the potential of the cell is:

"E = 0.34 - (-0.76) )- \frac{8.314\cdot298}{2\cdot96485.3399}\cdot\text{ln}\frac{0.5}{0.05}= 1.07" V.


Answer: the cell potential is 1.07 V.